a 25-mL volumetric flask with its stopper has a mass of 32.982g. the same flask, filled to the line with an ethylene glycol ( C2H6O2, automotive antifreeze) solution, has a mass of 58.198g. What is the density of the ethylene glycol solution (g/cm3)?
Answer #1

Flask Volume (V) - 2smi mass of Hask (m) = 329829 my a mass of Hask + S +6 %) = 58.1989 mass of G H O2 = Ma - m - 58.1989) -(